OpenStudy (anonymous):

Solve : 2 sin Ɵ = 1 please help !

OpenStudy (anonymous):

sin theta = 1/2 therefore sin^-1 (1/2) = 30 degrees

OpenStudy (cwrw238):

2 sin Ɵ = 1 sin Ɵ = 1/2 Ɵ = 30 , 150 degrees ( for Ɵ between 0 and 360 degrees)

OpenStudy (anonymous):

Thanks guys !

OpenStudy (cwrw238):

yw